OpenStudy (sarahc):

Let f(x) = 5x + 12. Find f–1(x).

OpenStudy (anonymous):

I was also confused on these.

OpenStudy (anonymous):

Switch y and x, y being f(x)

OpenStudy (anonymous):

is this question about finding the inverse of a function?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

f-1(x)

OpenStudy (anonymous):

f^-1

OpenStudy (anonymous):

\[y = 5x +12\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

x=5y+12

OpenStudy (anonymous):

\[x = 5y +12\]

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x-12=5y

OpenStudy (anonymous):

(x-12)/5=y

OpenStudy (anonymous):

\[5y = x -12\]

OpenStudy (anonymous):

\[y = \frac{ x-12 }{ 5 }\rightarrow f ^{-1}(x)=\frac{ x }{ 5 }-\frac{ 12 }{5 }\]

OpenStudy (anonymous):

Too late, I got it.

OpenStudy (anonymous):

I GTG.

OpenStudy (anonymous):

Good luck math.

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 5 }(x-12)\]

OpenStudy (sarahc):

so it would x-12/5?

OpenStudy (anonymous):

yup

OpenStudy (sarahc):

thanks!

OpenStudy (anonymous):

you are welcome but follow the steps

OpenStudy (anonymous):

so you can do it yourself

OpenStudy (anonymous):

if you practice the steps i showed you, you'll almost be able to solve any problem of the kind

OpenStudy (sarahc):

okay I did. I got 100 :)

OpenStudy (anonymous):

Good for you.

OpenStudy (anonymous):

did u get 100%

OpenStudy (anonymous):

good i am happy

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

sorry i was busy in the computer science section

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Anyway, can you help me?

OpenStudy (anonymous):

I posted my questions.

OpenStudy (anonymous):

sure what is ur qs? if i know how to solve it?

OpenStudy (anonymous):

Bumping now.

OpenStudy (anonymous):

so where is the question

OpenStudy (anonymous):

It should be the third one down.